OpenStudy (anonymous):

FREE MEDAL WITH ANSWER 2x + 1 / 3 = x + 6 / 2

OpenStudy (anonymous):

x=16

OpenStudy (anonymous):

2x+(1)/(3)=x+(6)/(2) Reduce the expression (6)/(2) by removing a factor of 2 from the numerator and denominator. 2x+(1)/(3)=x+3 Since x contains the variable to solve for, move it to the left-hand side of the equation by subtracting x from both sides. 2x+(1)/(3)-x=3 Since 2x and -x are like terms, add -x to 2x to get x. x+(1)/(3)=3 Since (1)/(3) does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting (1)/(3) from both sides. x=-(1)/(3)+3 Simplify the right-hand side of the equation. x=(8)/(3)

OpenStudy (anonymous):

@roza93 could you explain how you got that?

OpenStudy (anonymous):

@thewonderfuladele there were 2 other equations I had to answer and this problem had to have the same answer as at least one of the others, and 8/3 wasn't one of the previous answers. Only 15 and 16.

OpenStudy (anonymous):

\[\frac{ 2x+1}{ 3 } = \frac{ x+6 }{ 2 }cross multiply to get 2(2x+1)=3(x+6)\]

OpenStudy (anonymous):

4x+2 = 3x+18 x=16

OpenStudy (anonymous):

Thanks! @roza93